OpenStudy (anonymous):

Solve the equation x/2+1=square root 3 for booth exact irrational root and it's rational approximation to the nearest tenth

OpenStudy (anonymous):

\[ \frac{x}{2}+1=\sqrt{3}\\\frac{x}{2}=-1+\sqrt{3}\\x=-2+2\sqrt{3}\\x\approx 1.5 \]